### v1

# PatchRec: Multi-Grained Patching for Efficient LLM-based Sequential Recommendation

Abstract.
Large Language Models for sequential recommendation (LLM4SR), which transform user-item interactions into language modeling, have shown promising results. However, due to the limitations of context window size and the computational costs associated with Large Language Models (LLMs), current approaches primarily truncate user history by only considering the textual information of items from the most recent interactions in the input prompt. This truncation fails to fully capture the long-term behavioral patterns of users. To address this, we propose a multi-grained patching framework — PatchRec. It compresses the textual tokens of an item title into a compact item patch, and further compresses multiple item patches into a denser session patch, with earlier interactions being compressed to a greater degree. The framework consists of two stages: (1) Patch Pre-training, which familiarizes LLMs with item-level compression patterns, and (2) Patch Fine-tuning, which teaches LLMs to model sequences at multiple granularities. Through this simple yet effective approach, empirical results demonstrate that PatchRec outperforms existing methods, achieving significant performance gains with fewer tokens fed to the LLM. Specifically, PatchRec shows up to a 32% improvement in HR@20 on the Goodreads dataset over uncompressed baseline, while using only 7% of the tokens. This multi-grained sequence modeling paradigm, with an adjustable compression ratio, enables LLMs to be efficiently deployed in real-world recommendation systems that handle extremely long user behavior sequences.
Sequential Recommendation, Large Language Model, Multi-Grained Compression
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
(a) Truncating interaction sequences
(b) Performance of LLM4SR w.r.t. truncated sequence length
Figure 1. (a) Most LLM4SR works consider the latest-$K$ interactions in the complete user interaction sequence as the truncated historical item sequence feed to LLMs. (b) Performance of LLM-based recommender on Movielens-1M at different numbers of items in the truncated interaction sequence. The number of tokens in historical item sequence ranges from 39 to 552, excluding the tokens of task description.
Large Language Models (LLMs) (OpenAI, 2023; Google, 2023; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI, 2024) have demonstrated remarkable success across diverse domains, sparking a growing interest in their application to sequential recommendation (SR) (Wang et al., 2019a). A common paradigm for using LLMs as sequential recommender models (LLM4SR) (Geng et al., 2022; Yue et al., 2023; Zhang et al., 2023b; Bao et al., 2023b) involves two key steps: (1) formatting a user’s interaction history (i.e., a sequence of previously interacted items) into an input prompt suitable for LLMs, and (2) fine-tuning the LLMs to generate the target response (i.e., predicting the next item of interest). For the first step, most leading methods (Bao et al., 2023b; Liao et al., 2024b, a; Yang et al., 2023b) truncate each interaction history, retaining only the most recent $K$ items [^1] K, with common values being 5, 10, or 20, is typically much smaller than the total length of complete interaction history.. This truncation integrates a much shorter interaction sequence into the input prompt, primarily due to the context window size limitations and computational demand of LLMs. In such a truncated interaction sequence, each item is represented using the textual metadata (e.g., movie titles like “Gone with the Wind”) (Bao et al., 2023b, a; Liao et al., 2024a), as shown in Figure 1(a). While effective for capturing short-term preferences, this approach directly ignores interactions prior to the most recent $K$ items, making it challenging to capture the long-term interests encoded in extended item sequences. Modeling long sequences remains largely unexplored in LLM4SR, but the primary focus of this paper.
To investigate the impact of sequence length, we conduct preliminary experiments by supervised fine-tuning (SFT) Llama-3.2-1B-Instruct (Dubey et al., 2024) with different truncated sequence length on the Movielens dataset, as depicted in Figure 1(a). The results, illustrated in Figure 1(b), yield two key observations: (1) Performance improves with truncated sequence length: As the number of items in the truncated interaction sequence increases, recommendation accuracy — HR@20, improves substantially, showcasing the value of modeling longer interaction histories. For instance, earlier interactions with animated films contribute to accurately predicting subsequent items like “Alice Wonderland”. (2) Performance plateaus beyond a threshold: However, beyond a certain sequence length (e.g., 60 items), performance levels off with opaque increase. This plateau effect may result from the diminishing relevance of older interactions to a user’s current preferences and the potential introduction of noise (Zhou et al., 2018; Chen et al., 2022). Take Figure 1(a) as an example again, where older items like “Pulp Fiction” may serve as irrelevant or noisy signals in the user‘s interaction history.
These observations motivate us to consider the following challenges of modeling long sequences in LLM4SR:
  • Constraints of LLMs: The length of the item sequence significantly influences the input prompt length, which is constrained by the context window length of LLMs (Su et al., 2024) and the associated computational costs. How can we design approaches to efficiently model longer item sequences within these limitations to mine long-term user interest?
  • Temporal significance of interactions: Current LLM4SR methods do not explicitly regard the temporal distinctions of relevance (Ying et al., 2018), treating all items in a user’s interaction sequence as equally important (Liao et al., 2024b; Bao et al., 2023a). How can we incorporate mechanisms to capture the varying importance of interactions over time, thereby aligning more closely with evolving user preferences?
Figure 2. Performance-efficiency trade-off for vanilla LLM4SR method, PatchRec-I, and PatchRec-S on MovieLens-1M. PatchRec variants demonstrate improvements in both metrics compared to the baseline. The shaded region highlights the area above the performance-efficiency trade-off curve, representing desirable configurations.
To tackle these challenges, we introduce a hierarchical compression approach, as depicted in Figure 3. First, the textual tokens of item titles are compressed into compact item patches. These items are then organized into sessions, with the corresponding item patches further compressed into session patches, capturing higher-level patterns in user behavior. This compression allows for the accommodation of a significantly larger number of items within a fixed context window. Additionally, to account for the temporal significance of user interactions, we integrate multi-grained session patches, item patches, and raw textual tokens, enabling more compact and dynamic modeling of user interaction histories.
Building on this idea, we propose PatchRec — a new patching framework for LLM4SR. As illustrated in Figure 4, PatchRec operates in the following two stages:
  • Patch Pre-training. During each training step, the model simultaneously learns both the original uncompressed and compressed versions of the same data sample, thereby establishing a correspondence between the compressed item patches and their original textual tokens.
  • Patch Fine-tuning. This stage models the temporal variation in item importance using distinct compression granularities. Specifically, items from earlier interactions are compressed to a higher degree, while more recently interacted items are represented by a lower degree of compression, allowing the model to effectively capture longer user histories with varying levels of importance.
Extensive experiments are conducted on three benchmark datasets (i.e., MovieLens-1M (Harper and Konstan, 2016), Goodreads (Wan and McAuley, 2018), and MovieLens-100K (Harper and Konstan, 2016)), revealing the following key findings: (1) For the same truncated interaction sequence, PatchRec uses only 7.34% LLM input tokens compared to uncompressed SFT, while simultaneously improving HR@20 by up to 32% on the Goodreads dataset. (2) For the same computational cost, PatchRec models 3.44 times more user behaviors than uncompressed SFT, resulting in a substantial performance improvement of up to 13% on the MovieLens-1M dataset.
Our contributions can be summarized as follows:
  1. (1)
    We propose a multi-grained (i.e., both item- and session-level) compression framework — PatchRec, modeling users’ long-term historical behaviors in LLM4SR.
  2. (2)
    Our proposed pre-training and fine-tuning framework enables LLMs to internalize compressed representations as well as time-aware preferences for sequential modeling.
  3. (3)
    Experimental results demonstrate that our approach not only saves computational resources but also enhances recommendation performance.
2. Related Work
In this section, we provide a brief overview of existing works on LLM-based sequential recommendation (LLM4SR) and discuss user sequence modeling approaches in sequential recommendation.
2.1. LLM for Sequential Recommendation
Sequential recommendation involves predicting the next item a user is likely to interact with, based on their past engagements (Fang et al., 2020; Wang et al., 2019b). With LLMs demonstrating vast world knowledge and impressive reasoning capabilities across various domains (OpenAI, 2023; Google, 2023; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI, 2024; Shao et al., 2024), their potential for sequential recommendation tasks has garnered significant attention from researchers (Hou et al., 2024; Yue et al., 2023). Existing work on LLM4SR can be roughly classified into two categories: (1) LLMs as enhancers: In this approach, LLMs augment traditional recommender models by processing textual features associated with users and items (Yuan et al., 2023; Hou et al., 2022), and (2) LLMs as recommenders: LLMs directly generate recommendations for users’ next interactions (Gao et al., 2023; Bao et al., 2023b; Rajput et al., 2023).
The primary focus of this paper is on the latter — LLMs as recommenders — where the core challenge is converting user-item interaction sequences into a format suitable for language modeling task. Current approaches in LLM4SR can be broadly categorized into two types: (1) Text-based item representations: Items are represented using texts, such as item titles (Bao et al., 2023b, a), descriptions (Li et al., 2023), or index numbers (Geng et al., 2022), which are readily digestible for LLMs, and (2) Multi-modal token incorporation: Items are represented by tokens derived from other modalities (Liao et al., 2024b; Zhang et al., 2023a), such as embeddings projected from item representations learned by traditional recommender models (Tang and Wang, 2018; Hidasi et al., 2016a; Kang and McAuley, 2018).
In both aforementioned categories, each item is typically represented by multiple tokens. Due to the quadratic computational complexity $O(n^{2})$ of self-attention mechanism in LLMs w.r.t. sequence length  [^2] The computational complexity of vanilla attention mechanisms is $O(n^{2})$. Recent studies proposed methods to reduce it to approximately $O(n)$, using multiple tokens to represent one item in a user’s interaction sequence becomes computationally demanding. This motivates us to explore more efficient item and even sequence representations, aiming to improve both the performance and efficiency of LLM4SR methods.
2.2. User Sequence Modeling
Sequential recommendation interprets users’ interaction sequences as reflections of their underlying interests. Traditional sequential recommender models leverage neural networks to capture these patterns, employing architectures such as Recurrent Neural Networks (RNNs) (Hidasi et al., 2016a; Tan et al., 2016; Quadrana et al., 2017; Beutel et al., 2018; Hidasi et al., 2016b), Convolutional Neural Networks (CNNs) (Tang and Wang, 2018; Tuan and Phuong, 2017; Yuan et al., 2019), and attention mechanisms (Kang and McAuley, 2018; Sun et al., 2019). These models primarily focus on users’ latest-$K$ interactions, thus emphasizing short-term interests.
To incorporate information from richer historical behaviors, several studies have extended these approaches to model long-term interests. Methods include introducing static user representations (Zhang et al., 2018), leveraging two-layer hierarchical attention networks (Ying et al., 2018), periodically updating user memory (Ren et al., 2019), and employing multi-channel memory networks (Pi et al., 2019). To efficiently model long-term user behaviors, some works adopt linear attention mechanism (Liu et al., 2023) and mamba (Liu et al., 2024) in the architecture. Additionally, other works retrieve items from a user’s interaction history based on their relevance to the target item (Pi et al., 2020; Qin et al., 2022; Chen et al., 2021; Cao et al., 2022).
The integration of LLMs into recommendation systems introduces a unique challenge — further constraining the number of users’ historical behaviors that can be modeled due to the context window length of LLMs and the associated computational resource limitations. Despite this challenge, few studies have explored the modeling of long sequences in LLM4SR. One notable work (Lin et al., 2024) tackles the challenge by retrieving the top-$K$ interacted items from a user’s history based on semantic relevance to the target item. However, this retrieval-based approach is specifically designed for the click-through rate (CTR) prediction task and is not directly applicable to the next-item prediction task in sequential recommendation, which is the primary focus of this paper.
3. Method: PatchRec
Figure 3. Hierarchical compression. The textual tokens of an item title are aggregated into a compact item patch. Then several adjacent item patches are further compressed into a denser session patch.
Figure 4. Two-stage training framework of PatchRec. In the patch pre-training stage, we augment each uncompressed sequence $s$ with a compressed version $s^{\prime}$, where each item is independently and randomly compressed from raw textual tokens into an item patch, in order to build connections between item patches and textual item titles. During the training process, the degree of compression gradually increases from 0 to 1. In the patch finetuning stage, we fine-tune the LLMs with time-aware compressed sequences, allowing the LLM to become familiar with a mixed space of various compression granularities in downstream usage. For each sequence, interactions that occur earlier are compressed to a greater extent, while interactions that occur later are compressed to a lesser extent. PFT-S can achieve higher compression ratio with denser session patches.
We propose PatchRec, a simple yet effective framework designed to model longer historical item sequences in LLM-based sequential recommendation (LLM4SR) through hierarchical compression. This section begins with the task formation of LLM4SR, followed by an introduction to our hierarchical compression method. Finally, we detail the two-stage training process that underpins our framework.
3.1. Task Formation
Current LLMs, parameterized by $\Theta$, are typically trained using an autoregressive modeling approach, where the model generates each token $\mathbf{x}_{t}$ conditioned on the previous ones $\mathbf{x}_{<t}$ of the sequence. Formally, this is expressed as:
(1) $P(X;\Theta)=\prod_{t=1}^{T}P(\mathbf{x}_{t}|\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t-1};\Theta).$
In the context of sequential recommendation, LLM4SR utilizes LLMs as sequential recommenders to generate responses, i.e., predicting the next item that aligns with user preferences, based on an input prompt describing their historical interaction sequence. This task can be formalized as language modeling:
(2) $P(Y|X;\Theta)=\prod_{t=1}^{T}P(\mathbf{y}_{t}|\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N},\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{t-1};\Theta),$
where $X$ represents the LLM input sequence of $N$ tokens, $Y$ denotes the output sequence, and $\Theta$ are the model parameters. In this formulation, $X$ typically consists of a truncated interaction sequence of $K$ items accommodated in a prompt template. Each item title in $X$ is often represented by several textual tokens $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the hidden size of the LLM. The sequence $Y$ is intended to represent the predicted title of the next item.
3.2. Hierarchical Compression
To incorporate more interactions within truncated sequences, given the limitation of LLMs’ context window length and computation demand, we hierarchically compress the representation of historical item sequences, as shown in Figure 3.
3.2.1. Item Patch Compression.
In the context of LLM4SR, the textual tokens of each item title naturally form a unit that semantically represents the item. To compress these textual tokens into an item patch, we propose a simple yet effective method — computing the average of their token embeddings as follows:
(3) $\mathbf{z}^{i}_{j}=\frac{1}{|\mathcal{T}_{j}|}\sum_{t\in\mathcal{T}_{j}}\mathbf{x}_{t},$
where $\mathbf{z}^{i}_{j}\in\mathbb{R}^{d}$ denotes the item patch for the $j$-th item, $\mathcal{T}_{j}$ is the set of textual tokens associated with this item’s title, and $\mathbf{x}_{t}$ is the textual embedding of token $t$. As shown in Figure 3, the textual tokens of a movie title (e.g., “The Bridges of Madison Country”) are compressed into one compact item patch. This compression reduces the token-level granularity into a semantic-level representation of the item, effectively compressing the input space.
3.2.2. Session Patch Compression.
Since users often exhibit consistent interests over a period of time, consecutive interactions can be further compressed. To accomplish this, we partition a user’s historical item sequence into sessions, where each session groups together a predefined number of $L$ adjacent items  [^3] Alternative partitioning methods can also be employed, such as grouping items interacted with during a specific time window.. The item patches within each session are subsequently compressed into a session patch by averaging the embeddings of the item patches as follows:
(4) $\mathbf{z}^{s}_{j}=\frac{1}{|\mathcal{S}_{j}|}\sum_{t\in\mathcal{S}_{j}}\mathbf{z}^{i}_{t},$
where $\mathbf{z}^{s}_{j}\in\mathbb{R}^{d}$ denotes the session patch for the $j$-th session, $\mathcal{S}_{j}$ is the set of item patches associated with this session, and $\mathbf{z}^{i}_{t}$ represents the embedding of the $t$-th item patch. As depicted in Figure 3, every two adjacent item patches are further compressed into a denser session patch. This hierarchical compression strategy enables efficient handling of longer sequences while preserving the underlying user interest patterns.
3.3. Two-Stage Training
Although the compressed item/session patches $\mathbf{z}^{i}$/$\mathbf{z}^{s}$ maintain the same dimensionality as the original textual token, LLMs do not inherently understand them. To address this, we propose a two-stage training process, as illustrated in Figure 4. In the first stage, Patch Pre-training, we train the LLM to learn the compression patterns from textual tokens to item patches by comparing item patches with their corresponding original textual tokens. In the second stage, Patch Fine-tuning, we adapt the LLM to the final multi-grained and time-aware compressed format.
3.3.1. Patch Pre-training
For each interaction sequence in a training batch, as shown in Figure 4, where items are represented by their textual titles, we augment the interaction sequence with a compressed version of it. In this compressed interaction sequence, the textual tokens of each item are independently and randomly compressed into an item patch with probability $p$. The training objective is formulated as:
(5) $P(Y|X;\Theta)=P(Y|\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{k-1},\mathbf{z}^{i}_{j},\mathbf{x}_{k+n+1},\ldots,\mathbf{x}_{N};\Theta),$
where the $j$-th item’ title tokens $\mathbf{x}_{[k:k+n]}$ are compressed into an item patch $\mathbf{z}^{i}_{j}$ as defined in Equation 3. The probability $p$ gradually increases from 0 to 1 during training, following the schedule:
(6) $p=\frac{\tau}{T},$
where $\tau$ denotes the current training step, and $T$ represents the total number of training steps. By employing $p$ as a scheduler, the item representations in the compressed version progressively transition from textual tokens to item patches over the course of training.
Within each batch, interaction sequences in both raw and compressed formats are presented simultaneously. This approach establishes a clear correspondence for each interaction, effectively linking the textual tokens to their respective item patches.
3.3.2. Patch Fine-tuning
To incorporate the temporal decay of user interests in a sequence, we propose two fine-tuning strategies with different compression granularities: Patch Fine-tuning on Items (PFT-I) and Patch Fine-tuning on Sessions (PFT-S). As shown in Figue 4, they are designed to accommodate varying computational resource constraints.
[…]
(7) $P(Y|X;\Theta)=P(Y|\mathbf{z}^{i}_{1},\mathbf{z}^{i}_{2},\ldots,\mathbf{z}^{i}_{K-M},\mathbf{x}_{j},\ldots,\mathbf{x}_{N};\Theta),$
where $\mathbf{x}_{j}$ denotes the start textual token of the $(K-M+1)$-th item and $N$ is the total number of tokens across all $K$ truncated items. PFT-I achieves a moderate compression ratio.
PFT-S divides the truncated historical item sequence into groups of items based on interaction time, with each group containing at most $L$ adjacent items. The interactions in the latest group are retained as original textual tokens, while those in the second-latest group are compressed into item patches. All earlier groups are further compressed into the densest session patches. The training objective is formulated as:
(8) $P(Y|X;\Theta)=P(Y|\mathbf{z}^{s}_{1},\mathbf{z}^{s}_{2},\ldots,\mathbf{z}^{i}_{1},\mathbf{z}^{i}_{2},\ldots,\mathbf{x}_{1},\mathbf{x}_{2},\ldots,;\Theta).$
PFT-S achieves a higher compression ratio.
Fine-tuning the LLM on these multi-grained compressed interaction sequences enables it to effectively capture time-aware user interest patterns, striking a balance between representation effectiveness and computational efficiency.
4. Experiments
In this section, we present a comprehensive comparison of PatchRec against various baseline methods, including traditional recommender models, efficient long-term recommender models, and LLM-based recommenders, across three real-world datasets: MovieLens-1M, Goodreads, and MovieLens-100K. Our objective is to demonstrate the effectiveness and efficiency of PatchRec under different levels and granularities of compression. Additionally, we analyze the mechanisms through which our approach achieves its superior performance. Furthermore, we analyze the performance of PatchRec in user groups with more long-term behaviors.
4.1. Experimental Settings
Table 1. Statistics of datasets. Int.: Interactions. TT.: Title tokens.
Dataset # User # Item # Int. # Max Int. # Avg. Int. # Avg. TT.
MovieLens-1M 6,037 3,883 575,281 1,435 95.28 3.94
Goodreads 2,090 8,793 162,539 430 77.77 9.82
MovieLens-100K 682 1,682 51,824 378 75.99 3.86
4.1.1. Datasets
We conduct extensive experiments on three real-world benchmark datasets: MovieLens-1M, Goodreads, and MovieLens-100K. The MovieLens datasets are widely used benchmarks, sourced from a movie recommendation platform  [^5] https://movielens.org/, while the Goodreads dataset is derived from the largest online community for readers and book recommendations  [^6] https://www.goodreads.com/.
To ensure data quality, we apply filtering based on rating thresholds of 3 and 5 for MovieLens and Goodreads, respectively. For MovieLens-100K, we retain only users with more than 20 interactions. For Goodreads, we exclude users who have interacted with fewer than 50 items and items that have received fewer than 10 interactions from distinct users. To prevent information leakage and balance computational overhead during training and evaluation, we partition the interactions in each dataset into training, validation, and test splits, following a temporal split ratio of 48:1:1 based on timestamps. The dataset statistics are summarized in Table 1.
4.1.2. Implementations
We adopt Llama-3.2-1B-Instruct (Dubey et al., 2024) as the backbone LLM for our framework, fine-tuning all its parameters. Each training stage consists of one epoch and requires approximately 12 NVIDIA A800 GPU hours. To balance computational efficiency and resource allocation, we use a training batch size of 64 for MovieLens datasets and 32 for Goodreads, and a test batch size of 16 for all three datasets. The training employs a cosine learning rate scheduler with a warmup ratio of 0.05, and the learning rates are set to 8e-6,5e-5,1e-5 for MovieLens-1M, Goodreads, and MovieLens-100K, respectively. For each interaction, the latest $K$ historical interactions are selected as the truncated user sequence. All experiments access item titles that appear within their respective datasets to ensure fair and consistent evaluation.
Table 2. Performance comparison with 100 items in the truncated user sequence. CR is short for compression ratio. Bold and underlined indicate the best and the second-best results, respectively. PatchRec improves recommendation accuracy with less computational demands. *(p-value $\ll$ 0.05)
MovieLens-1M* Goodreads* MovieLens-100K
Model HR@10 N@10 HR@20 N@20 CR HR@10 N@10 HR@20 N@20 CR HR@10 N@10 HR@20 N@20 CR
GRU4Rec 0.0623 0.0298 0.1106 0.0419 - 0.0401 0.0219 0.0592 0.0267 - 0.0661 0.0307 0.1211 0.0445 -
Caser 0.0384 0.0180 0.0712 0.0262 - 0.0116 0.0057 0.0220 0.0083 - 0.0501 0.0234 0.0910 0.0336 -
SASRec 0.0594 0.0286 0.1035 0.0397 - 0.0404 0.0229 0.0569 0.0271 - 0.0605 0.0290 0.1069 0.0406 -
LinRec 0.0622 0.0296 0.1063 0.0407 - 0.0251 0.0125 0.0398 0.0161 - 0.0619 0.0283 0.1048 0.0393 -
Mamba4Rec 0.0648 0.0303 0.1116 0.0421 - 0.0281 0.0146 0.0450 0.0188 - 0.0667 0.0309 0.1086 0.0415 -
MoRec 0.0341 0.0164 0.0616 0.0233 - 0.0122 0.0061 0.0236 0.0089 - 0.0383 0.0184 0.0762 0.0279 -
Vanilla 0.0933 0.0396 0.1482 0.0491 1.00 0.0563 0.0236 0.0770 0.0249 1.00 0.0986 0.0443 0.1680 0.0611 1.00
PatchRec-I 0.1058 0.0455 0.1616 0.0525 3.44 0.0733 0.0289 0.0976 0.0293 6.81 0.0967 0.0416 0.1738 0.0574 3.38
PatchRec-S 0.0967 0.0408 0.1526 0.0496 9.23 0.0748 0.0295 0.1013 0.0301 13.62 0.1016 0.0450 0.1680 0.0554 3.85
4.1.3. Evaluation
To evaluate recommendation performance, we use two key metrics: HitRatio@10/20 and NDCG@10/20. Additionally, to assess efficiency, we introduce the Compression Ratio, defined as the ratio of the number of tokens in the historical item sequence before compression to the number after compression. A higher compression ratio indicates greater efficiency.
4.2. Main Performance Comparison
4.2.1. Baselines
We denote the models trained using patch pre-training and PFT-I as PatchRec-I, and those trained with patch pre-training and PFT-S as PatchRec-S. Our method is compared against a diverse set of baselines, including traditional recommender models, efficient long-term sequential recommender models, and LLM-based recommenders.
  • Traditional Recommender Models. We select three sequential recommender models as representatives of different mechanisms for capturing user behavior patterns: GRU4Rec(Hidasi et al., 2016a) (RNN-based), Caser(Tang and Wang, 2018) (CNN-based) and SASRec(Kang and McAuley, 2018) (attention-based)  [^7] These models are implemented following (Yang et al., 2023a)..
  • Efficient Long-term Recommender Models. LinRec (Liu et al., 2023) and Mamba4Rec (Liu et al., 2024) are chosen as representatives of efficient long-term sequential recommendation methods. LinRec leverages linear attention mechanisms, while Mamba4Rec employs Mamba blocks (Gu and Dao, 2023) for efficient modeling of long-term user behaviors.
  • LLM-based Recommenders. We include three representative methods that incorporate LLMs in sequential recommendation: (1) MoRec  [^8] We adopt BERT as the encoder for items’ textual metadata and SASRec as the recommender backbone. (Yuan et al., 2023) enhances traditional recommenders by incorporating textual features encoded by LLMs. (2) Vanilla LLM4SR (Vanilla) is adapted from leading methods such as TALLRec (Bao et al., 2023b) and BIGRec (Bao et al., 2023a) by applying SFT on truncated interaction sequences without compression, while maintaining evaluation settings consistent with PatchRec. It serves as the most important control group for PatchRec, with the only difference being whether the historical item sequence is compressed.
(a) MovieLens-1M
(b) Goodreads
(c) MovieLens-100K
Figure 5. Performance comparison between PatchRec-I with vanilla with the same item numbers in interaction sequence. The compression ratios of PatchRec-I are 2.27, 3.06, and 2.25 for MovieLens-1M, Goodreads, and MoviLens-100K, respectively. Impressively, with improved efficiency, PatchRec-I does not compromise recommendation accuracy.
(a) MovieLens-1M
(b) Goodreads
(c) MovieLens-100K
Figure 6. Performance comparison between the PatchRec-I with vanilla with comparable numbers of tokens in the interaction sequence. The item number in the interaction sequence is 100 for PatchRec-I, with various compression ratios (cr). For PatchRec-I, the latest-$M$ ($M$=3, 5, 10, 20, 30, 50, 100) items are represented with their textual titles, and earlier items are compressed into item patches. PatchRec-I delivers superior recommendation performance compared to vanilla at the same computational cost.
4.2.2. Results
The truncation lengths of all the methods compared in Table 2 are 100 (i.e., modeling the latest-100 items for each sequence). For PatchRec-I, we keep the latest-$M$ ($M$=5) interactions as items’ textual titles and compress the early interactions into item patches. For PatchRec-S, we use a group size $L$ of 5 for MovieLens-1M and Goodreads, and 20 for MovieLens-100K; so the latest-$L$ interactions are retained as textual tokens, the $L$ interactions in the second latest group are compressed into item patches, and the earlier interactions are compressed into session patches.
The results in Table 2 demonstrate that both PatchRec-I and PatchRec-S outperform all traditional recommender, efficient long-term sequential recommender, and LLM-based recommender baselines across all four metrics on MovieLens-1M and GoodReads. And on MovieLens-100K, with comparable recommendation accuracy, PatchRec achieves a much short input sequence.
Notably, PatchRec-I achieves HR@20 improvements of 9.04%, 26.75%, and 3.45% over vanilla, with compression ratios of 3.44, 6.81, and 3.38 on MovieLens-1M, Goodreads, and MovieLens-100K, respectively. Similarly, PatchRec-S achieves HR@20 improvements of 2.97% and 31.56% on MovieLens-1M and Goodreads, respectively, with higher compression ratios of 9.23 and 13.62. On MovieLens-100K, PatchRec-S achieves comparable performance to vanilla with a compression ratio of 3.85.
These results indicate that both PatchRec-I and PatchRec-S can significantly enhance recommendation performance while drastically reducing the number of LLM input tokens. Overall, the experimental findings highlight the effectiveness of our approach in simultaneously improving recommendation accuracy and reducing computational overhead.
4.3. Analysis of PatchRec-I
To further assess the performance of PatchRec-I in details, we compare it with vanilla under two distinct experimental settings across three datasets, addressing the following research questions:
  • RQ1: Does compression result in a loss of recommendation performance when processing identical interaction sequences?
  • RQ2: For a fixed inference computational cost, does compression improve recommendation performance?
4.3.1. Performance Comparison under the Same Items in the Interaction Sequence (RQ1)
We maintain the same truncation length for interaction sequences, with item counts of 20, 40, 60, 80, and 100, and compare HR@20 of PatchRec-I with vanilla LLM4SR to examine whether the compression of early items in PatchRec-I leads to any loss of recommendation performance. The results are shown in Figure 5.
On the MovieLens-1M and Goodreads datasets, PatchRec-I consistently outperforms vanilla, with relative gains exceeding 5% and 30%, respectively. This improvement may be attributed to the increased information density in the compressed item patches, which mitigates the influence of noisy interactions. On the MovieLens-100K dataset, PatchRec-I achieves performance comparable to vanilla, indicating that the proposed compression strategy does not compromise recommendation accuracy.
Furthermore, under a fixed compression ratio, we observe that as the truncation length (i.e., the items in the interaction sequence) increases across the three datasets, the recommendation performance of PatchRec-I generally improves, unlike vanilla, which doesn’t show steady improvement. This trend suggests that, for PatchRec-I, incorporating more historical interactions enhances the ability to model long-term user behaviors effectively.
(a) Goodreads
(b) MovieLens-100K
Figure 7. The performance-efficiency trade-off of vanilla, PatchRec-I, and PatchRec-S on Goodreads and MovieLens-100K dataset. In the scatter plot, points that are positioned further towards the upper right corner indicate better performance and higher efficiency. both PatchRec-I and PatchRec-S provide a better performance-efficiency trade-off compared to vanilla.
4.3.2. Performance Comparison under Comparable Numbers of Tokens (RQ2)
We compare the recommendation performance of PatchRec-I with vanilla with comparable token counts for LLM input to evaluate the impact of compression on recommendation performance at an equivalent computational cost. The results are shown in Figure 6.
In all experiments across the three datasets, PatchRec-I processes 100 items, which are represented as item patches or raw textual titles. However, due to varying compression ratios, the number of tokens required differs. For each bar in the figures, both PatchRec-I and vanilla input the same number of tokens. Despite this parity in token usage, PatchRec-I consistently achieves higher HR@20 than vanilla, with relative improvements of at least 7%, 20%, and 2%, and reaching up to 13%, 72%, and 19% on MovieLens-1M, Goodreads, and MovieLens-100K, respectively.
These results demonstrate that across all three datasets, PatchRec-I delivers superior recommendation performance compared to vanilla at the same computational cost, underscoring the efficiency and effectiveness of the proposed compression method.
4.4. Analysis of PatchRec-S
Table 3. The performance comparison of PatchRec-I/S between with and without patch pre-training on MovieLens-1M dataset. The patch pre-training stage improves recommendation performance.
Method HR@10 N@10 HR@20 N@20
PatchRec-I
w/o Patch Pre-training 0.1014 0.0439 0.1550 0.0504
w/ Patch Pre-training 0.1058 0.0455 0.1616 0.0525
PatchRec-S
w/o Patch Pre-training 0.0920 0.0386 0.1436 0.0463
w/ Patch Pre-training 0.0967 0.0408 0.1526 0.0496
PatchRec-S further compresses the earliest item patches into session patches, thereby achieving even higher compression ratio. To validate its effectiveness, we compare the compression ratio and HR@20 of three methods — vanilla, PatchRec-I, and PatchRec-S — while maintaining the same number of 100 items in the interaction sequence across different compression ratios on three datasets. The results are shown in Figure 2 and Figure 7.
The figures show that the distribution of PatchRec-I and PatchRec-S points is positioned further towards the upper-right corner compared to vanilla across all three datasets. Specifically:
  • PatchRec-S consistently outperforms vanilla in HR@20 across various compression ratios on the MovieLens-1M and Goodreads datasets. Even at compression ratios as high as 9.23 and 13.62 for MovieLens-1M and Goodreads, respectively, PatchRec-S maintains superior HR@20 compared to vanilla.
  • PatchRec-I points are generally positioned higher on the performance axis, while PatchRec-S points are shifted further to the right, reflecting greater efficiency.
These results indicate that both PatchRec-I and PatchRec-S provide a better performance-efficiency trade-off compared to vanilla. Furthermore, the distinction between the two methods demonstrates their complementary strengths: PatchRec-I prioritizes performance, while PatchRec-S emphasizes efficiency.
4.5. Effectiveness of the Patch Pre-training Stage
Table 4. Performance comparison between vanilla, patch pre-training, and data augmentation with dropout on MovieLens-1M dataset. The enhancements achieved by compression are not simply due to the random exclusion of items.
Method HR@10 N@10 HR@20 N@20
Vanilla 0.0933 0.0396 0.1482 0.0491
Dropout 0.0936 0.0402 0.1480 0.0492
Patch Pre-training 0.1026 0.0431 0.1610 0.0532
In this section, we ablate to demonstrate the benefits of the patch pre-training stage.
4.5.1. Impact of the Patch Pre-training Stage
To evaluate the importance of the patch pre-training stage, we compare the recommendation performance of PatchRec-I and PatchRec-S under two settings: (1) Starting from LLM checkpoints trained with the patch pre-training stage, which is denoted as w/ patch pre-training. (2) Directly initialized from the Llama-3.2-1B-Instruct checkpoint without patch pre-training, denoted as w/o patch pre-training. The evaluation is conducted on the MovieLens-1M dataset, and the results are shown in Table 3. From the table, we can observe that both PatchRec-I and PatchRec-S with the patch pre-training stage consistently outperforms those without that stage.
This demonstrates the critical role of the patch pre-training stage in improving performance. During this stage, the inclusion of both uncompressed textual tokens (i.e., item titles) and compressed item patches within the same batch enables the LLM to effectively learn the patterns of item-level compression. This adaptation allows the LLM to transition seamlessly from the textual token space to the compressed item patch space, laying a robust foundation for the subsequent multi-grained patch fine-tuning stage.
4.5.2. On Data Augmentation
To examine the performance improvements attributed to the compression operations in the patch pre-training stage, we compare our compression approach against a baseline method that directly drops the items compressed during the patch pre-training data augmentation process.
The experimental results, as summarized in Table 4, reveal that patch pre-training consistently outperforms both the vanilla (i.e., no-compression) and dropout baselines. This demonstrates that the enhancements achieved by our compression strategy are not simply due to the random exclusion of items. Our compression design retains critical patterns while increasing the effective information density of the input. These findings highlight the strength of our method in balancing efficiency and effectiveness, ensuring that compression enhances.
4.6. Performance and Efficiency for Long-term User Behaviors
Figure 8. Performance (i.e., HR@20) and efficiency (i.e., inference time) comparison of vanilla, PatchRec-I, and PatchRec-S for long sequences (i.e., users with at least 500 behaviors) on MovieLens-1M dataset. PatchRec enhances both performance and efficiency in modeling long-term user behaviors.
To further evaluate the performance and efficiency of PatchRec when applied to even longer sequences, we analyze user groups with more than 500 interactions in MovieLens-1M. This experiment was conducted using a Nvidia A100 GPU with 40 GB of memory. For each test interaction, we maximize the number of historical items visible to all three methods (i.e., vanilla LLM4SR, PatchRec-I, and PatchRec-S) to fully utilize the available GPU memory, which simulates the pragmatism in real-world recommender systems that make full use of all available computing sources and data. Specifically, we truncate the Latest-$K$ ($K$=150, 350) historical items for both vanilla and PatchRec-I, whereas PatchRec-S is capable of accommodating all historical items (1435 at most) within the 40 GB GPU memory constraints.
In the case of PatchRec-I, the latest 20 interactions are retained as textual titles, while earlier interactions are compressed into item patches. For PatchRec-S, the parameter $L$, which denotes the number of items in a session, is set to 20. The experimental results are shown in Figure 8. The figure demonstrates that PatchRec-I achieved a 25.11% improvement in HR@20 with only 12.99% inference time, compared to vanilla. Similarly, PatchRec-S achieved a 20.31% increase in HR@20 relative to vanilla, with a even higher 13.09-fold reduction in inference time. These results indicate that our proposed methods significantly enhance both performance and efficiency in modeling long-term user behaviors, implying that our PatchRec is capable of modeling practical historical behavior lengths via LLM4SR.
5. Future Work
While PatchRec demonstrates promising computational efficiency and recommendation performance, certain limitations remain, which we aim to address in future research:
  • Advanced Compression Designs. Although our proposed compression method is simple yet effective and more suitable for multi-grained patch alignment, there is significant potential for exploring more sophisticated designs. Future work may investigate advanced text-to-item and item-to-session compression methods beyond simple averaging, as well as innovative grouping techniques (i.e., consolidating all items within a specific time window into a single session).
  • Adaptive Multi-Grained Compression. Currently, we employ a uniform, predefined multi-grained compression strategy across all sequences. A promising avenue for future exploration involves developing adaptive approaches that dynamically apply multi-grained compression based on available computational resources. This could include determining, for each sequence, the optimal number of items to compress into session patches, item patches, or retain as raw textual titles.
  • Lifelong Sequence Modeling. While PatchRec enhances efficiency in modeling long-term user behaviors with LLMs, there remains an open challenge in effectively modeling lifelong user histories or infinitely long interaction sequences within constrained computational budgets. Future efforts could focus on innovative architectures or methodologies to address this issue.
By addressing these directions, we aim to further enhance the scalability, flexibility, and effectiveness of PatchRec in real-world recommendation systems.
6. Conclusion
In this paper, we introduce PatchRec, a simple yet effective compression framework designed for multi-grained modeling of users’ extensive historical behaviors in LLM-based sequential recommendation (LLM4SR). To address the significant challenges posed by LLM constraints and the temporal dynamics of user interactions — mostly untouched in prior works — PatchRec leverages a hierarchical and time-aware compression strategy, enabling a deeper and more nuanced understanding of user preferences. By adopting a two-stage training process, PatchRec effectively adapts LLMs to operate within a multi-grained sequence representation space. Extensive experiments on benchmark datasets demonstrate that PatchRec not only dramatically reduces the input token length for LLMs but also consistently improves recommendation performance.
This work paves the way for future research into more efficient and effective ways to harness the power of LLMs in sequential recommendation. We hope PatchRec will inspire further exploration into efficient LLM-based sequential recommenders for lifelong user sequence modeling.